OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{Solve }\hspace{.33em}\\~\\ 3x^2+2y^2=5xy~~and~~ 12x+12y=5xy \hspace{.33em}\\~\\ x,y\in \mathbb{R} }\end{align}\)

OpenStudy (freckles):

12x+12y=5xy 12x=5xy-12y 12x=y(5x-12) y=12x/(5x-12) So I would try to replace the y's in the first equation with 12x/(5x-12) and solve for x

OpenStudy (mathmath333):

painful

OpenStudy (freckles):

you think we can find a pretty trick?

OpenStudy (freckles):

like it is one of those questions?

OpenStudy (mathmath333):

yes i hope

OpenStudy (freckles):

well like I don't know if this helps but we have an ellispe if we write 3x^2+2y^2=12x+12y

OpenStudy (mathmath333):

if it were only integers not real numbers then how could we do it

OpenStudy (mathmath333):

wolfram does gives two pairs of integers http://www.wolframalpha.com/input/?i=solve+3x%5E2%2B2y%5E2%3D5xy%2C12x%2B12y%3D5xy+over+reals

ganeshie8 (ganeshie8):

\(3x^2+2y^2=5xy\) factors like below \[(x-y)(3x-2y)=0\] we get two cases : when \(x =y\), secod equation becomes \[12x+12x=5x\cdot x \implies x = 0, ~\frac{24}{5}=y \] you an work the other case similarly..

ganeshie8 (ganeshie8):

fun fact : \(ax^2+by^2+cxy=0\) represents a pair of straight lines passing through origin, so we can always factor it as below product : \[ax^2+by^2+cxy =0= (y-m_1x)(y-m_2x)\]

OpenStudy (mathmath333):

i will rote this fact